OpenStudy (anonymous):

Find a number that is 24 more than its opposite.

Parth (parthkohli):

Opposite?

OpenStudy (lgbasallote):

hint: the opposite of 24 is its negative

Parth (parthkohli):

So ambiguous questions these days. lol

OpenStudy (anonymous):

So is the equation is 24x=-24?

OpenStudy (lgbasallote):

no..

Parth (parthkohli):

No... Hint: If your number is ♥, then its opposite is -♥ and your equation is ♥ = 24 + (-♥)

OpenStudy (lgbasallote):

i said opposite of 24 is -24...in your case it's opposite of a number (not 24)

OpenStudy (anonymous):

So in the equation there's no positive 24

OpenStudy (lgbasallote):

there is... positive 24 would be the "24 more than" more than means add so 24 more than means add 24 make sense?

Parth (parthkohli):

Yes, @abb50 is right.

OpenStudy (anonymous):

yes it does

OpenStudy (anonymous):

Let x be the number. Since it is 24 more than its opposite, then x - (-x) = 24. This means 2x = 24 or x = ???

OpenStudy (anonymous):

opposite may mean 42 .

OpenStudy (anonymous):

No, it should be -24

OpenStudy (anonymous):

a=-24+24=0